OpenStudy (samigupta8):

Lim (x-sinx/x) sin(1/x) When x----->0 ...

OpenStudy (samigupta8):

The division sign is under the whole term x-sinx and not just sinx

OpenStudy (anonymous):

is it this question!

OpenStudy (samigupta8):

@faiqraees

OpenStudy (samigupta8):

Nope!... It is sin(1/x) this term is in multiplication with first one.. And even x tends to 0

hartnn (hartnn):

you're stuck only fir sin(1/x) term right? what is the range of sine function?

OpenStudy (samigupta8):

Exactly! (-1,1)

OpenStudy (samigupta8):

Oops...closed bracket i meant...

hartnn (hartnn):

ok, so whatever number x tends to sin (1/x) will ALWAYS be withing that finite range. agree?

hartnn (hartnn):

within*

OpenStudy (samigupta8):

It will be oscillating at an indeed rapid pace between the values -1,1

OpenStudy (anonymous):

it doesn't exist. and it diverges.

hartnn (hartnn):

right, we'll take your words :) what is this product now: 0*(number oscillating at an indeed rapid pace between the values -1,1) = ...?

OpenStudy (samigupta8):

Bt this is not exact 0...the term u multiplied it with

hartnn (hartnn):

the 2nd term? its a finite number right?

OpenStudy (samigupta8):

Nope...the first one

hartnn (hartnn):

what? 1- sin x/x = 1-1 = 0 basics of limits :P

OpenStudy (anonymous):

is it this!

OpenStudy (samigupta8):

Sinx/x <1

OpenStudy (anonymous):

can you write the actual question which make sense .. put bracket to distinguish!

hartnn (hartnn):

ok, let me put it this way: \(\lim \limits_{x \to 0}\dfrac{\sin x}{x}=1\) agree?

OpenStudy (samigupta8):

Yep...

OpenStudy (samigupta8):

Do we need to separate limits like this Lim (f(x)*g(x)) can be written as Lim (f(x)) *lim(g(x))

hartnn (hartnn):

exactly

OpenStudy (samigupta8):

Ok fine then...it"s done...

OpenStudy (samigupta8):

Thanks...

OpenStudy (zarkon):

you need to use the squeeze theorem

OpenStudy (samigupta8):

Thanks sir...! Bt it's an mcq type question